OpenStudy (anonymous):

Please help!!! Use implicit differentiation to find dy/dy of x^3+xy=y^2-1

OpenStudy (anonymous):

dy/dy?...

OpenStudy (anonymous):

\[\large x^3+xy=y^2-1\] IMPLICIT DIFFERENTIATION: \[\large D_{x}(x^3+xy)=D_{x}(y^2-1)\] \[\large 3x^2+y+x(\frac{dy}{dx})=2y(\frac{dy}{dx})\] \[\large 3x^2+y=\frac{dy}{dx}(2y-x)\] \[\large \frac{dy}{dx}=\frac{3x^2+y}{2y-x}\]